a 1000 kg truck is parked on a bridge 100 m long at a distance of 25 m from one end of the bridge. the mass of the bridge is 200 kg m. calculate the reaction forces at the supports at the two ends of the bridge. (take g 10 ms )
1
Expert's answer
2019-03-22T10:41:10-0400

The equilibrium conditions are as follows: R2L = mga + 0.5MgL, R1L = mg (L-a) + 0.5MgL and R1 + R2 = mg + Mg, from which follows R2 = mga/L + 0.5MgL, R1 = mg (1-a/L) + 0.5Mg, where M = ρL, therefore, R1 = 10 (250 + 100*100) = 102500 (N), R2 = 10 (750 + 100*100) = 107500 (N).


Finally, R1 = 102500 N, R2 = 107500 N.